# Jailbreaking Leading Safety-Aligned LLMs with Simple Adaptive Attacks

### v2

@@ 4 Jailbreaking Leading Safety-Aligned LLMs
In this section, we detail the adaptive attacks we have developed for several families of leading safety-aligned LLMs. We provide a detailed descriptions of the main evaluations here and show the rest in Table 22 in the appendix where we also present results on Vicuna-13B, Mistral-7B, Phi-3-Mini, and Nemotron-4-340B.
@@ 4.1 Jailbreaking Llama-2, Llama-3, and Gemma Models
Here, we focus on open-weights Llama-2-Chat (7B, 13B, 70B parameters) (Touvron et al., 2023), Llama-3-Instruct (8B parameters) and Gemma-7B models (Google, 2023). These models have been significantly safety-trained, rendering them resilient to jailbreaks even in white-box scenarios (Zou et al., 2023).
Approach.
The key element to jailbreak the Llama models is self-transfer, where successful adversarial suffixes found by RS on simpler requests are used as initialization for RS on more complex requests. Notably, these adversarial strings tend to be to some extent transferable across different model sizes (e.g., from 7B to 13B models), but for the best result we repeat the self-transfer procedure for each model size separately. The same approach is also successful on Gemma-7B, although prompt + RS alone already demonstrates high attack success rate.
Results.
Table 3 shows that we achieve 100% attack success rate on all these models. For Llama-2-Chat models, our standard adversarial prompt templates yield a 0% success rate, confirming the effectiveness of their safety alignment. When we apply Prompt + RS the attack success rate (ASR) increases to 48%. Ultimately, our composite attack strategy—which combines prompting, random search, and self-transfer—achieves a 100% attack success rate for all LLMs, surpassing all existing methods. For Llama-2-Chat-7B, the best reported success rate is 92% by PAP (Zeng et al., 2024) which is an LLM-assisted method. However, this method requires 10 restarts to achieve such accuracy, and its success rate drops to 46% with only one restart. In contrast, for this model, one restart is sufficient for our method to achieve a 100% ASR. Meanwhile, for the 13B and 70B models, Mazeika et al. (2024) reports ASR below 40%, while there is no prior evaluation available for Llama-3-Instruct and Gemma-7B since they are relatively recent models.
Table 3: Llama-2, Llama-3, and Gemma models. We report the attack success rate using the GPT-4 judge.
@@ Results.
Llama-2-Chat-7B TAP Zeng et al. (2024) 4%
Llama-2-Chat-7B PAIR Chao et al. (2023) 10%
Llama-2-Chat-7B GCG Chao et al. (2023) 54%
Llama-2-Chat-7B PAP Zeng et al. (2024) 92%
@@
Llama-2-Chat-13B TAP Mazeika et al. (2024) 14%*
Llama-2-Chat-13B PAIR Mazeika et al. (2024) 15%*
Llama-2-Chat-13B GCG Mazeika et al. (2024) 30%*
@@
Llama-2-Chat-70B TAP Mazeika et al. (2024) 13%*
Llama-2-Chat-70B PAIR Mazeika et al. (2024) 15%*
Llama-2-Chat-70B GCG Mazeika et al. (2024) 38%*
@@
Convergence plots.
We show convergence curves in Figure 2, where we plot the average logprob of the token ‘Sure’ and average attack success rate for three representative models (Llama-3-Instruct-8B, Llama-2-Chat-7B, and Gemma-7B) for random search (RS) with and without self-transfer. The plot confirms that starting from a good initialization via self-transfer is key for query efficiency and high success rate.
Figure 2: Convergence curves. We show the average logprob of the token ’Sure’ and attack success rate for three representative models (Llama-3-Instruct-8B, Llama-2-Chat-7B, and Gemma-7B) with and without self-transfer. Starting from a good initialization via self-transfer is key for query efficiency and high attack success rate.
4.2 Jailbreaking R2D2 Model
Approach.
Similarly to Llama-2-Chat, the standard prompt template, alone or with RS, shows limited effectiveness. However, in contrast to Llama-2-Chat, self-transfer is ineffective here. Motivated by the fact that R2D2 was trained to refuse a specific prompt structure (i.e., request and adversarial suffix), we circumvent safety guardrails by using a different prompt structure. We use an in-context learning prompt (see Table 8 in the appendix), which we found the model to be particularly sensitive to. We use random search on top of the in-context prompt to maximize the probability of the initial token “Step” (instead of “Sure”) to be consistent with the new prompt template.
Results.
As shown in Table 4, using the in-context prompt alone achieves a 90% attack success rate, which RS boosts to 100%. This significantly surpasses the 61% reported by Mazeika et al. (2024) using TAP (Mehrotra et al., 2023). Interestingly, the in-context prompt is less effective on other models like Llama-2-Chat (see Table 22 in the appendix).
Table 4: R2D2 from HarmBench. We report the attack success rate according to the GPT-4 judge.
Model Method Source Success rate
R2D2-7B GCG Mazeika et al. (2024) 6%
R2D2-7B PAIR Mazeika et al. (2024) 48%
R2D2-7B TAP Mazeika et al. (2024) 61%
@@ 4.3 Jailbreaking GPT Models
GPT models are the most popular state-of-the-art LLMs with non-trivial built-in safety features. We use the API checkpoints gpt-3.5-turbo-1106 and gpt-4-1106-preview for our experiments.
Approach.
We observed that GPT-3.5 Turbo is extremely brittle to manually designed prompts, with no need for more sophisticated techniques. In contrast, GPT-4 Turbo demonstrates greater resistance to these adversarial prompt templates. Thus, for this model, we rely on self-transfer to achieve more successful jailbreaks.
Results.
Table 5 summarizes our results: with the prompt template alone, we achieve 100% success rate on GPT-3.5 Turbo, outperforming the baselines. For GPT-4 Turbo, using the prompt alone leads only to 28% success rate. However, by combining the prompt, RS, and self-transfer, we improve the previous best ASR from 59% (Mazeika et al., 2024) to 96%. For reference, we also provide baselines with standard GPT-4 (i.e., not Turbo) in Table 5 but we do not evaluate it ourselves due to its higher costs.
@@ Results.
GPT-3.5 Turbo PAIR Chao et al. (2023) 60%
GPT-3.5 Turbo TAP Zeng et al. (2024) 80%
GPT-3.5 Turbo GCG Zeng et al. (2024) 86%
GPT-3.5 Turbo PAP Zeng et al. (2024) 94%
@@
GPT-4 PAP Zeng et al. (2024) 92%
GPT-4 Turbo PAIR Mazeika et al. (2024) 33%*
GPT-4 Turbo TAP Mazeika et al. (2024) 36%*
GPT-4 Turbo TAP-Transfer Mazeika et al. (2024) 59%*
@@
Non-determinism in GPT-3.5/4.
The limitation of the API providing only the top-5 log-probabilities is not critical, as it is often straightforward to prompt a desired token, like “Sure”, to appear in the top-5. A more challenging issue is the non-deterministic output, since RS does not necessarily have a correct signal to refine the adversarial string. As illustrated in Figure 3, identical queries can yield varying log-probabilities, even with a fixed seed parameter and temperature zero in the API. The randomness makes random search less effective, although it still succeeds to a large extent.
@@
Figure 3: Non-determinism of GPT models. The histogram of log-probabilities for the first response token using the same query repeated 1 00010001\,0001 000 times for GPT-4 Turbo. We use temperature zero and we fix the seed parameter in the API, but the returned log-probabilities are still non-deterministic.
@@ 4.4 Jailbreaking Claude Models
Claude models are known for their high safety levels. In line with this, Anthropic does not provide access to logprobs for these models which prevents direct iterative attack like random search. Thus, we first test a transfer attack using an adversarial suffix optimized on GPT-4 with random search. We enhance the attack with multiple random restarts to leverage different generations with temperature one. Then we investigate an attack method that utilizes Anthropic’s prefilling feature, [^1] https://docs.anthropic.com/claude/docs/prefill-claudes-response which is not commonly available from other LLM API providers like OpenAI. To improve the attack success rate, we use the prefilling feature together with our prompt (Table 2) which we split into a system and user part (see Tables 17 and 18 for a detailed ablation).
Transfer attack.
As shown in Table 6, the direct transfer attack is particularly effective on certain models such as Claude 3 Sonnet (100% ASR). Given Claude-3’s recent release in early March 2024, there are no established baselines for comparison. The attack success rate of the transfer attack improves when the initial segment of the prompt (which corresponds to the set of rules to follow) is provided as the system prompt. In this way, we can achieve 100% ASR on Claude 2.0 and 98% ASR on Claude 3 Haiku. We present an illustrative example of a transfer attack on Claude 3 Sonnet in Figure 1 and postpone more complete results to the appendix (Table 16). We conclude that while Claude models exhibit increased robustness against static harmful requests, their resistance to adversarial suffixes—challenging to derive without logprobs—is not perfect.
@@
Claude 2.0 PAP Zeng et al. (2024) 0%
Claude 2.0 GCG Chao et al. (2023) 4%
Claude 2.0 PAIR Chao et al. (2023) 4%
@@
Claude 2.0 Transfer from GPT-4 + system prompt Ours 100%
Claude 2.0 Prefilling attack Ours 100%
@@
Claude 2.1 Transfer from GPT-4 Ours 0%
Claude 2.1 Prefilling attack Ours 100%
Claude 3 Haiku Transfer from GPT-4 + system prompt Ours 98%
Claude 3 Haiku Prefilling attack Ours 100%
Claude 3 Sonnet Transfer from GPT-4 Ours 100%
Claude 3 Sonnet Prefilling attack Ours 100%
Claude 3 Opus Transfer from GPT-4 Ours 0%
Claude 3 Opus Prefilling attack Ours 100%
@@
Prefilling attack.
The prefilling feature makes jailbreaking straightforward on Claude models, even without any search (Table 6). For comparison, the previous best result on Claude 2.0 is 61% (Shah et al., 2023) while we get 100% using only up to 10 random restarts. The latest Claude 2.1 model, released in November 2023, is significantly more robust to both transfer and prefilling attacks. Nonetheless, we are able to get $100\%$ ASR with 100 restarts. We note that GPT-4 as a semantic judge sometimes has false positives, more often so on Claude 2.1. At the same time, the technical report of Claude-3 (Anthropic, 2024) mentions fewer refusals in its release announcement, which, in our findings, correlated with vulnerability to jailbreaking. We provide more complete experimental results, including the number of restarts in each case, in Tables 17 and 18 in the appendix.
@@ 5 Adaptive Attacks for Trojan Detection
Setup.
Rando & Tramèr (2023) showed the possibility of implanting backdoor attacks during the RLHF training of LLMs by poisoning a small percentage of the preference data with a universal suffix. Then a model that typically refuses to answer harmful queries can then be jailbroken by appending the suffix to any request. Rando & Tramèr (2024) recently launched a competition to retrieve backdoor attacks in five Llama-2-7B models, each poisoned with a different trojan. A reward model was also provided to evaluate the safety of prompt-response pairs (higher scores to safer responses), alongside a dataset of harmful requests. The objective is to discover triggers (5 to 15 tokens long) acting as universal jailbreaks for each model.
Approach.
Random search could be directly applied to optimize the score provided by the reward model on some training examples. However, despite the triggers being relatively short, the search space is extremely large, as the vocabulary $T$ of the Llama-2 tokenizer comprises 32001 tokens, and straightforward RS becomes particularly inefficient. It is noteworthy that the five LLMs, denoted by $M_{1},\ldots,M_{5}$, were fine-tuned from the same base model, thereby sharing the weights initialization, including those of the embedding matrix that maps tokens to the LLM’s continuous feature space (each token $t_{i}$ is associated with a vector $v_{i}\in\mathbb{R}^{4096}$, for $i=0,\ldots,32000$). Given that the tokens part of the trigger appear abnormally frequently, we anticipate that their corresponding embedding vectors significantly deviate from their initial values. Building on this intuition, for any pair of models $M_{r}$ and $M_{s}$ with embedding matrices $v^{r}$ and $v^{s}$, we compute the distance $\left\|v^{r}_{i}-v^{s}_{i}\right\|_{2}$ for each token, sorting them in decreasing order $\pi^{rs}$, where
@@
The final pool of candidate trigger tokens for a model $M_{r}$ is the intersection of such sets: $\textrm{cand}(M_{r})=\bigcap_{s\neq r}\textrm{top-}k(M_{r},M_{s})$. Given that the five models are fine-tuned using different random subsets of the training data, this approach is approximate but narrows down the candidate tokens to a manageable pool (e.g., $k=1000$ yields $|\textrm{cand}(M_{r})|\in[33,62]$ for $r=2,\ldots,5$, $|\textrm{cand}(M_{1})|=480$), which makes random search feasible. Our strategy to identify jailbreaking triggers for the poisoned model $M_{r}$ involves conducting a random search in the token space over the set $\textrm{cand}(M_{r})$. We restrict the search to triggers of five tokens, as this length yielded the best results. In each iteration, we filter out candidate triggers that do not start with a blank space, contain blank spaces or are not invariant to decoding-encoding, [^2] Given a sequence of token indices, the tokenizer decodes it into a text string. However, re-encoding this string via the tokenizer does not guarantee the recovery of the initial sequence. following the competition hints. The objective minimized by RS is the average score of the reward model on a batch of training examples, aiming to ensure the trigger’s universality (generalization to unseen prompts).
@@
true trojans -11.96 -7.20 -5.76 -4.93 -7.63 -37.48
Results.
In Table 7 we report the average scores of the reward model over a held-out test set of harmful prompts for the five models, and their sum: without the triggers, the models produce safe answers (high scores), indicating proper alignment. We then compare the effectiveness of the triggers discovered by competing methods (those ranked 2nd and 3rd in the competition) with our approach: RS on the restricted set of tokens achieves the best (lowest) score for 3 out of 5 target models, as well as the best overall score. Moreover, the scores achieved by our method are not far from those given by the exact trojans, i.e. used to poison the datasets. We note that the numbers from Table 7 marginally differ from the ones reported in Rando & Tramèr (2024): first, our top-1 entry was slightly better due to the usage of gradient guidance for some models. Second, we re-evaluated all solutions, as well as the models without triggers and with the true trojans, on the evaluation set [^3] The first half of the dataset is available at https://huggingface.co/datasets/ethz-spylab/competition_eval_dataset. which led to slightly different values for all methods, but same ranking. To conclude, similarly to our approach for jailbreaking, our method includes an adaptive component (the selection of candidate token pools) that leverages task-specific information, complemented by an automated optimization process through RS.
6 Discussion, Recommendations, Outlook, and Limitations
Discussion.
Our work makes a few important methodological contributions that we would like to further highlight and discuss.

### v3

@@ 4 Jailbreaking Leading Safety-Aligned LLMs
In this section, we detail the adaptive attacks we have developed for several families of leading safety-aligned LLMs, i.e., we focus primarily on production-grade models. We provide a detailed descriptions of the main evaluations here and show the rest in Table 21 in the appendix, where we also present results on Vicuna-13B, Mistral-7B, Phi-3-Mini, and Nemotron-4-340B.
@@ 4.1 Jailbreaking Llama-2, Llama-3, and Gemma Models
Here, we focus on open-weight Llama-2-Chat (7B, 13B, 70B parameters) (Touvron et al., 2023), Llama-3-Instruct-8B, and Gemma-7B models (Google, 2023). For the Llama models, we use the safety prompt from Touvron et al. (2023) (see Section A.1 for more details) that improves resilience to jailbreaks. These models have undergone significant safety training, rendering them resilient to jailbreaks, even in white-box scenarios (Zou et al., 2023).
Approach. The key element to jailbreak the Llama models is self-transfer, where successful adversarial suffixes found by random search on simpler requests are used as initialization for random search on more complex requests. Notably, these adversarial strings tend to be to some extent transferable across different model sizes (e.g., from 7B to 13B models), but for the best result we repeat the self-transfer procedure for each model size separately. The same approach is also successful on Gemma-7B, although prompt + random search alone already demonstrates high attack success rate.
Results. Table 2 shows that we achieve 100% attack success rate on all these models. For Llama-2-Chat models, our standard adversarial prompt templates yield a 0% success rate, confirming the effectiveness of their safety alignment. When we apply Prompt + random search the attack success rate (ASR) increases to 48%. Ultimately, our composite attack strategy—which combines prompting, random search, and self-transfer—achieves a 100% attack success rate for all LLMs, surpassing all existing methods. For Llama-2-Chat-7B, the best reported success rate is 92% by PAP (Zeng et al., 2024) which is an LLM-assisted method. However, this method requires 10 restarts to achieve such accuracy, and its success rate drops to 46% with only one restart. In contrast, for this model, one restart is sufficient for our method to achieve a 100% ASR. Meanwhile, for the 13B and 70B models, Mazeika et al. (2024) reports ASR below 40%, while there is no prior evaluation available for Llama-3-Instruct and Gemma-7B since they are relatively recent models.
Table 2: Llama, Gemma, and R2D2. We report the attack success rate using the GPT-4 judge.
@@
Llama-2-Chat-7B Tree of Attacks with Pruning Zeng et al. (2024) 4%
Llama-2-Chat-7B Prompt Automatic Iterative Refinement Chao et al. (2023) 10%
Llama-2-Chat-7B Greedy Coordinate Gradient Chao et al. (2023) 54%
Llama-2-Chat-7B Persuasive Adversarial Prompts Zeng et al. (2024) 92%
@@
Llama-2-Chat-13B Tree of Attacks with Pruning Mazeika et al. (2024) 14%*
Llama-2-Chat-13B Prompt Automatic Iterative Refinement Mazeika et al. (2024) 15%*
Llama-2-Chat-13B Greedy Coordinate Gradient Mazeika et al. (2024) 30%*
@@
Llama-2-Chat-70B Tree of Attacks with Pruning Mazeika et al. (2024) 13%*
Llama-2-Chat-70B Prompt Automatic Iterative Refinement Mazeika et al. (2024) 15%*
Llama-2-Chat-70B Greedy Coordinate Gradient Mazeika et al. (2024) 38%*
@@
R2D2-7B Greedy Coordinate Gradient Mazeika et al. (2024) 6%
R2D2-7B Prompt Automatic Iterative Refinement Mazeika et al. (2024) 48%
R2D2-7B Tree of Attacks with Pruning Mazeika et al. (2024) 61%
@@
Convergence plots. We show convergence curves in Figure 2, where we plot the average logprob of the token ‘Sure’ and average ASR for three representative models (Llama-3-Instruct-8B, Llama-2-Chat-7B, and Gemma-7B) for random search (RS) with and without self-transfer. The plot confirms that starting from a good initialization via self-transfer is key for query efficiency and a high ASR.
Figure 2: Convergence curves. We show the average logprob of the token “Sure” and attack success rate for three representative models (Llama-3-Instruct-8B, Llama-2-Chat-7B, and Gemma-7B) with and without self-transfer. Starting from a good initialization via self-transfer is key for query efficiency and high attack success rate.
4.2 Jailbreaking R2D2 Model
Approach. Similarly to Llama-2-Chat, the standard prompt template, alone or with RS, shows limited effectiveness. However, in contrast to Llama-2-Chat, self-transfer is ineffective here. Motivated by the fact that R2D2 was trained to refuse a specific prompt structure (i.e., request and adversarial suffix), we circumvent safety guardrails by using a different prompt structure. We use an in-context learning prompt (see Table 6 in the appendix), which we found the model to be particularly sensitive to. We use random search on top of the in-context prompt to maximize the probability of the initial token “Step” (instead of “Sure”) to be consistent with the new prompt template.
Results. As shown in Table 2, using the in-context prompt alone achieves a 90% attack success rate, which random search boosts to 100%. This significantly surpasses the 61% reported by Mazeika et al. (2024) using TAP (Mehrotra et al., 2023). Interestingly, the in-context prompt is less effective on other models like Llama-2-Chat (see Table 21 in the appendix).
@@ 4.3 Jailbreaking GPT Models
GPT models remain the most popular LLMs and have non-trivial safety alignment. We consider the following API checkpoints: gpt-3.5-turbo-1106, gpt-4-1106-preview, and gpt-4o-2024-05-13.
Approach. Since December 2023, OpenAI has made the predicted probabilities of their models available via the API, which we leverage for random search. We observed that GPT-3.5 Turbo is extremely brittle to manually designed prompts, with no need for more sophisticated techniques. In contrast, GPT-4 Turbo and GPT-4o demonstrate greater resistance to these adversarial prompt templates. Thus, for these models, we rely on random search and self-transfer to achieve more successful jailbreaks. Additionally, for GPT-4o, we customize the prompt and split it onto a system and user parts, see Table 7 in Appendix A.1 for details.
Results. Table 3 summarizes our results: with the prompt template alone, we achieve 100% success rate on GPT-3.5 Turbo, outperforming the baselines. For GPT-4 Turbo, using the prompt alone leads only to 28% success rate. However, by combining the prompt, RS, and self-transfer, we improve the previous best ASR from 59% (Mazeika et al., 2024) to 96%. Our default prompt template is completely ineffective against GPT-4o, resulting in 0% ASR. Moreover, running random search on top of it does not lead to a significantly higher ASR. However, our custom template (see Appendix A.1) already leads to 72% ASR, and we get to 100% ASR with random search and self-transfer. This large difference in the ASR again shows the importance of adaptive attacks that are customized for a particular model.
@@
GPT-3.5 Turbo Prompt Automatic Iterative Refinement Chao et al. (2023) 60%
GPT-3.5 Turbo Tree of Attacks with Pruning Zeng et al. (2024) 80%
GPT-3.5 Turbo Greedy Coordinate Gradient Zeng et al. (2024) 86%
GPT-3.5 Turbo Persuasive Adversarial Prompts Zeng et al. (2024) 94%
@@
GPT-4 Turbo Prompt Automatic Iterative Refinement Mazeika et al. (2024) 33%*
GPT-4 Turbo Tree of Attacks with Pruning Mazeika et al. (2024) 36%*
GPT-4 Turbo Tree of Attacks with Pruning (Transfer) Mazeika et al. (2024) 59%*
@@
GPT-4o Prompt Ours 0%
GPT-4o Custom Prompt Ours 72%
GPT-4o Custom Prompt + Random Search + Self-Transfer Ours 100%
@@
Non-determinism in GPT-3.5/4. The limitation of the API providing only the top-20 log-probabilities is not critical, as it is often straightforward to prompt a desired token, such as “Sure”, to appear in the top-20. A more challenging issue is the non-deterministic output, since random search does not necessarily have a correct signal to refine the adversarial string. Identical queries can yield varying log-probabilities, as shown in Figure 3, even with a fixed seed and temperature zero in the API. The randomness makes random search less effective, although it still succeeds to a large extent.
@@
Figure 3: Non-determinism of GPT models. The histogram of log-probabilities for the first response token using the same query repeated 1 00010001\,0001 000 times for GPT-4 Turbo. We use temperature zero and we fix the seed parameter in the API, but the returned log-probs are still non-deterministic.
@@ 4.4 Jailbreaking Claude Models
Claude models are known for their high safety levels. In line with this, Anthropic does not provide access to logprobs for these models which prevents direct iterative attack like random search. Thus, we first test a transfer attack using an adversarial suffix optimized on GPT-4 with random search. We enhance the attack with multiple random restarts to leverage different generations with temperature one. Then we investigate an attack method that utilizes Anthropic’s prefilling feature (Anthropic, 2024a), which is not commonly available from other LLM API providers like OpenAI. To improve the attack success rate, we use the prefilling feature together with our prompt (Figure 1) which we split into a system and user part (see Tables 16 and 17 for a detailed ablation).
Transfer attack. As shown in Table 4, the direct transfer attack is especially effective on specific models, such as Claude 3 Haiku, 3 Sonnet, and 3.5 Sonnet, with ASR rates of 98%, 100%, and 96%, respectively. Given the recent release of Claude 3 and Claude 3.5, there are no established baselines for comparison. The attack success rate of the transfer attack improves when the initial segment of the prompt (which corresponds to the set of rules to follow) is provided as the system prompt. In this way, we can achieve 100% ASR on Claude 2.0 and 98% ASR on Claude 3 Haiku. We present an illustrative example of a transfer attack on Claude 3 Sonnet in Figure 5 and show more complete results in the appendix (Table 15). We conclude that while Claude models exhibit increased robustness against static harmful requests, their resistance to adversarial suffixes—challenging to derive without logprobs—is not perfect.
Prefilling attack. The prefilling feature, combined with our prompt template, makes jailbreaking with 100% ASR straightforward on all Claude models—including Claude 3 and 3.5—even without any search, as we show in Table 4. For comparison, the previous best result on Claude 2.0 is 61% (Shah et al., 2023) while we get 100% using only up to 10 random restarts. The Claude 2.1 model appears to be the most robust model in the Claude series and is significantly more robust to both transfer and prefilling attacks. Although we are able to get $100\%$ ASR with 100 restarts, we note that GPT-4, as a semantic judge, sometimes produces false positives, particularly for this model. We provide more complete experimental results, including the number of restarts in each case, in Tables 16 and 17 in the appendix.
@@
Claude 2.0 Persuasive Adversarial Prompts Zeng et al. (2024) 0%
Claude 2.0 Greedy Coordinate Gradient Chao et al. (2023) 4%
Claude 2.0 Prompt Automatic Iterative Refinement Chao et al. (2023) 4%
@@
Claude 2.0 Prompt + Transfer from GPT-4 Ours 100%
Claude 2.0 Prompt + Prefilling Attack Ours 100%
@@
Claude 2.1 Prompt + Transfer from GPT-4 Ours 0%
Claude 2.1 Prompt + Prefilling Attack Ours 100%
Claude 3 Haiku Prompt + Transfer from GPT-4 Ours 98%
Claude 3 Haiku Prompt + Prefilling Attack Ours 100%
Claude 3 Sonnet Prompt + Transfer from GPT-4 Ours 100%
Claude 3 Sonnet Prompt + Prefilling Attack Ours 100%
Claude 3 Opus Prompt + Transfer from GPT-4 Ours 0%
Claude 3 Opus Prompt + Prefilling Attack Ours 100%
Claude 3.5 Sonnet Prompt + Transfer from GPT-4 Ours 96%
Claude 3.5 Sonnet Prompt + Prefilling Attack Ours 100%
@@ 5 Adaptive Attacks for Trojan Detection
Setup. Rando & Tramèr (2023) showed the possibility of implanting backdoor attacks during the RLHF training of LLMs by poisoning a small percentage of the preference data with a universal suffix. Then a model that typically refuses to answer harmful queries can then be jailbroken by appending the suffix to any request. Rando & Tramèr (2024) recently launched a competition to retrieve backdoor attacks in five Llama-2-7B models, each poisoned with a different trojan. A reward model was also provided to evaluate the safety of prompt-response pairs (higher scores to safer responses), alongside a dataset of harmful requests. The objective is to discover triggers (5 to 15 tokens long) acting as universal jailbreaks for each model.
Approach. Random search could be directly applied to optimize the score provided by the reward model on some training examples. However, despite the triggers being relatively short, the search space is extremely large, as the vocabulary $T$ of the Llama-2 tokenizer comprises 32001 tokens, and straightforward random search becomes particularly inefficient. It is noteworthy that the five LLMs, denoted by $M_{1},\ldots,M_{5}$, were fine-tuned from the same base model, thereby sharing the weights initialization, including those of the embedding matrix that maps tokens to the LLM’s continuous feature space (each token $t_{i}$ is associated with a vector $v_{i}\in\mathbb{R}^{4096}$, for $i=0,\ldots,32000$). Given that the tokens part of the trigger appear abnormally frequently, we anticipate that their corresponding embedding vectors significantly deviate from their initial values. Building on this intuition, for any pair of models $M_{r}$ and $M_{s}$ with embedding matrices $v^{r}$ and $v^{s}$, we compute the distance $\left\|v^{r}_{i}-v^{s}_{i}\right\|_{2}$ for each token, sorting them in decreasing order $\pi^{rs}$, where
@@
The final pool of candidate trigger tokens for a model $M_{r}$ is the intersection of such sets: $\textrm{cand}(M_{r})=\bigcap_{s\neq r}\textrm{top-}k(M_{r},M_{s})$. Given that the five models are fine-tuned using different random subsets of the training data, this approach is approximate but narrows down the candidate tokens to a manageable pool (e.g., $k=1000$ yields $|\textrm{cand}(M_{r})|\in[33,62]$ for $r=2,\ldots,5$, $|\textrm{cand}(M_{1})|=480$), which makes random search feasible. Our strategy to identify jailbreaking triggers for the poisoned model $M_{r}$ involves conducting a random search in the token space over the set $\textrm{cand}(M_{r})$. We restrict the search to triggers of five tokens, as this length yielded the best results. In each iteration, we filter out candidate triggers that do not start with a blank space, contain blank spaces or are not invariant to decoding-encoding, [^2] Given a sequence of token indices, the tokenizer decodes it into a text string. However, re-encoding this string via the tokenizer does not guarantee the recovery of the initial sequence. following the competition hints. The objective minimized by random search is the average score of the reward model on a batch of training examples, aiming to ensure the trigger’s universality (generalization to unseen prompts).
@@
Ground truth trojans -11.96 -7.20 -5.76 -4.93 -7.63 -37.48
Results. In Table 5 we report the average scores of the reward model over a held-out test set of harmful prompts for the five models, and their sum: without the triggers, the models produce safe answers (high scores), indicating proper alignment. We then compare the effectiveness of the triggers discovered by competing methods (those ranked 2nd and 3rd in the competition) with our approach: RS on the restricted set of tokens achieves the best (lowest) score for 3 out of 5 target models, as well as the best overall score. Moreover, the scores achieved by our method are not far from those given by the exact trojans, i.e. used to poison the datasets. We note that the numbers from Table 5 marginally differ from the ones reported in Rando & Tramèr (2024): first, our top-1 entry was slightly better due to the usage of gradient guidance for some models. Second, we re-evaluated all solutions, as well as the models without triggers and with the true trojans, on the evaluation set [^3] The first half of the dataset is available at https://huggingface.co/datasets/ethz-spylab/competition_eval_dataset. which led to slightly different values for all methods, but same ranking. To conclude, similarly to our approach for jailbreaking, our method includes an adaptive component (the selection of candidate token pools) that leverages task-specific information, complemented by an automated optimization process through random search.
6 Discussion, Recommendations, and Limitations
Our findings. Our work makes a few important methodological contributions.